Q1. Dart Company developed the following standard costs for its product for 2016:

DART COMPANY

Standard Cost Card                                                     

Cost Elements Standard Quantity   * Standard Price     = Standard Cost
Direct materials      4 pounds $5 $20
Direct labor       2 hours 10 20
Variable overhead     2 hours 4 8
Fixed overhead   2 hours 2 4
$52

                                                                                                                                                                                                              The company expected to work at the 120,000 direct labor hours level of activity and produce 60,000 units of product.

Actual results for 2016 were as follows:

  • 56,800 units of product were actually produced.
  • Direct labor costs were $1,092,000 for 112,000 direct labor hours actually worked.
  • Actual direct materials purchased and used during the year cost $1,108,800 for 231,000 pounds.

    Required:

    Compute the following variances for Dart Company for 2016 and indicate whether the variance is favorable or unfavorable. (show calculation!)

    1.     Total direct materials variance=____________________________________________________________________

    2.     Direct materials price variance=____________________________________________________________________

    3.     Direct materials quantity variance=____________________________________________________________________

    4.     Total direct labor variance=____________________________________________________________________

    5.     Direct labor price variance=____________________________________________________________________

    6.     Direct labor quantity variance=____________________________________________________________________
Answer #1
Variance analysis is a technical jargon to explain a situation where actual result differs from planned
or expected results. It is an act of comparing standards with actual. The study of variances help in
decision making and finding out the major areas which needs immediate attention.
Dart Company
Variances
Total DM variance = Standard cost - Actual cost
( 56800 * 20 ) - 1108800 = $ 27200 ( fav. )
MPV = ( SP - AP ) * AQ purchased
MPV = ( 5 - 4.8 ) * 231000 = $ 46200 ( favourable ) ( AP = 1108800 / 231000 )
MQV = ( SQ - AQ ) SP
MQV = ( 56800 * 4 - 231000 ) * 5 = $ 19000 ( unfavourable )
Total DL variance = Standard cost - Actual cost
( 56800 * 20) - 1092000 = $ 44000 ( fav. )
LRV = ( SR - AR ) AH
LRV = ( 10 - 9.75 ) 112000 = $ 28000 ( Favourable) ( AR = 1092000 / 112000 )
LEV = ( SH - AH ) SR
LEV = ( 56800 * 2 - 112000 ) 10 = $ 16000 ( favourable)